OpenStudy (anonymous):

A swimming is 18.4 ft in diameter and 4.4 ft deep. Determine the volume, V, of the pool in gallons if one gallon is 231cubic inches. ???? :(

OpenStudy (anonymous):

when you say it is 1.4ft in diameter, i suppose the swimming pool is in the shape of a sphere?

OpenStudy (anonymous):

Hmm upon closer inspection, it is circular pool, not a sphere. my mistake. ' Volume of the pool = area of circle * depth =pi * (18.4/2)^2 * 4.4 =1170 cubic ft, Given 1 gallon is 231 cubic inches... One cubic feet is 1728 cubic inch. Total cubic inches = 1170*1728= 2 021 760 cubic inches Number of gallons= 8752.2 gallons